OpenStudy (anonymous):

Find the product: (x + 2y)^3

OpenStudy (unklerhaukus):

\[(x+2y)^3=(x+2y)\times(x+2y)\times(x+2y)\]

OpenStudy (unklerhaukus):

\[=(x^2+4xy+4y^2)(x+2y)\]

OpenStudy (unklerhaukus):

\[=x^3+2x^2y+4x^2y+8xy^2+4xy^2+8y^3\] \[=x^3+6x^2y+12xy^2+8y^3\]

OpenStudy (anonymous):

These are the possible answers: A. x^3 + 8y^3 B. x^3 + 6xy + 8y^3 C. x^3 + 12x^2y + 6xy^2 + 8y^3 D. x^3 + 6x^2y + 12xy^2 + 8y^3 Is it D?

OpenStudy (anonymous):

Is it D @UnkleRhaukus ?

OpenStudy (unklerhaukus):

\[=x^3 + 6x^2y + 12xy^2 + 8y^3\]is exactly the same as D. x^3 + 6x^2y + 12xy^2 + 8y^3

OpenStudy (anonymous):

Oh okkay thanks! : )